What is the ratio of the velocity of a 3.53-MeV β ray to that of an α particle with the same kinetic energy? This should confirm that βs travel much faster than αs even when relativity is taken into consideration.
Answer #1

3.53 = (r - 1) x 0.511

r = 7.90

Vb = 3 x 10^8 x sqrt (1 - (1/7.9)^2) = 2.97 x 10^8 m/s

For alpha,

3.53 = (r - 1) x 4.0026 x 931.5

r = 1.00094

Va = sqrt(1 - (1/r))^2 = 1.30 x 10^7 m/s

Vb/Va = 22.77

Comment in case any doubt please rate my answer....